C₆H₁₂O₆+6O₂→6CO₂+6H₂O+energy. Consider respiration, one of the most common chemical reactions on earth. What mass of carbon dioxide forms in the reaction of 25 grams of glucose with 40 grams of oxygen? Ar; C=12.0, H=1.0, O=16.0


1
Expert's answer
2021-09-20T02:26:57-0400

for oxygen we have 6(16"\times 2" )=192

number of mole of oxygen="40\/192=0.2083"

mole ratio; "6CO_2:6O_2=6:6=1:1"

RMM of "CO_2=12+16\times2=48"

mass of carbon monoxide"=2083\times48 =9.9984g"